# Exchangeability and randomness for infinite and finite sequences

### v1

@@ Abstract
Randomness (in the sense of being generated in an IID fashion) and exchangeability are standard assumptions in nonparametric statistics and machine learning, and relations between them have been a popular topic of research. This note draws the reader’s attention to the fact that, while for infinite sequences of observations the two assumptions are almost indistinguishable, the difference between them becomes very significant for finite sequences of a given length.
The version of this note at http://alrw.net (Working Paper 45) is updated most often.
@@ 1 Introduction
In this note we will discuss two important assumptions about sequences of observations, exchangeability and randomness, using the word ‘‘randomness’’ in a somewhat old-fashioned sense of the individual observations being independent and identically distributed (following [19, Chap. 7], which used the standard terminology for its time, and [27]).
The relationship between exchangeability and randomness is very different in the cases of infinite sequences and finite sequences. In the former case, there is hardly any difference between the two assumptions. But in the latter, the difference may be vast. The study of this relationship has a long history, which will also be briefly reviewed in this note.
@@
According to Kolmogorov, and it is difficult to argue with his point of view, infinite sequences are empirically vacuous; we never observe them in reality. Therefore, he always insisted on either studying finite sequences or at least keeping them in mind. In fact, de Finetti’s results can also be stated in terms of finite sequences, provided we consider sequences of different lengths. In this note we will briefly discuss two modes of statistical hypothesis testing involving finite sequences, online and batch; the former corresponds to de Finetti’s setting and the latter to Kolmogorov’s.
@@
This note has little novelty, and all of its mathematical results are very simple. But the reader might find the scale of the difference between exchangeability and randomness for infinite $\mathbf{Z}$ surprising.
@@ 2 De Finetti’s theorem
Suppose we observe a data sequence $z_{1},z_{2},\dots\in\mathbf{Z}$ consisting of observations $z_{n}\in\mathbf{Z}$ that are elements of some measurable space, the observation space. The assumption of randomness is standard in machine learning: the data is coming from the product probability measure $Q^{\infty}$ for some $Q\in\mathfrak{P}(\mathbf{Z})$, where $\mathfrak{P}(\mathbf{Z})$ stands for the measurable space of all probability measures on $\mathbf{Z}$ (with the $\sigma$-algebra generated by the mappings $Q\mapsto Q(A)$, $A$ being an event in $\mathbf{Z}$).
Remark 2.1.
Effects of various distribution shifts have also been widely studied, but in this note we concentrate on the basic IID case. In machine learning we often have observations $z=(x,y)$ consisting of an object $x$ and its label $y$, but we do not insist on this.
@@
By de Finetti’s theorem each exchangeable probability measure $R$ on $\mathbb{R}^{\infty}$ is a mixture of product distributions:
@@
for some $\mu\in\mathfrak{P}(\mathfrak{P}(\mathbf{Z}))$. This was established by de Finetti [4, Chap. 4]. Hewitt and Savage [11, Theorem 7.3 and its discussion later in Sect. 7] notice that we can trivially replace $\mathbb{R}$ by any Borel space $\mathbf{Z}$ and point out that, to the best of their knowledge, every measurable set known to have importance in applied science is Borel. For example, every Polish space (complete separable metric space with its Borel $\sigma$-algebra) is Borel. On the other hand, there exists a separable metric space with its Borel $\sigma$-algebra for which (1) can be violated [8]. Let us assume in the rest of this note that the observation space $\mathbf{Z}$ is Borel.
It is well known that de Finetti’s theorem fails if we simply replace $\infty$ by a finite $N$ in its statement; see, e.g., [3, Sect. 4.7.1] (and [6, Sect. 1] for a further discussion of the extent to which de Finetti’s theorem can fail for $N=2$ and $\mathbf{Z}=\{0,1\}$).
@@
$\Gamma(z_{1},\dots,z_{n}):=\bigl\{(z_{n+1},z_{n+2},\dots):(z_{1},z_{2},\dots,z_{n+1},z_{n+2},\dots)\notin A\bigr\}$ (5)
for the future observations after observing $z_{1},\dots,z_{n}$, where $A$ is a critical region at some significance level $\epsilon$. Under both exchangeability and randomness, the coverage probability of this prediction set will be at least $1-\epsilon$.
There are two popular generalizations of critical regions, p-variables and e-variables, and both also produce identical results under exchangeability and randomness. Let us first check this for p-variables. According to the general definition (see, e.g., [21, Definition 1.2]), an exchangeability p-variable is a random variable $P:\mathbf{Z}^{\infty}\to[0,1]$ such that, for all $\epsilon\in(0,1)$,
@@
Since $\operatorname{\mathbb{P}^{\mathrm{X}}}=\operatorname{\mathbb{P}^{\mathrm{R}}}$, exchangeability and randomness p-variables coincide.
Again according to the general definition [21, Definition 1.2] (see also [10, (1)]), an exchangeability e-variable is a measurable function $F:\mathbf{Z}^{\infty}\to[0,\infty]$ such that $\operatorname{\mathbb{E}^{\mathrm{X}}}(F)\leq 1$, where
@@
3 Moving on to finite sequences of observations
@@ 3.1 Batch setting
$\frac{N}{N}\frac{N-1}{N}\dots\frac{1}{N}=\frac{N!}{N^{N}}\sim\sqrt{2\pi N}e^{-N};$ (11)
@@
Remark 3.1.
@@
Similarly to (5), we can invert exchangeability and randomness critical regions in the batch mode and use them for one-step-ahead prediction (whereas prediction in (5) was infinitely many steps ahead). Given an observed sequence $z_{1},\dots,z_{n}$, we can output
@@
as our prediction set for the next observation, where $A$ is a critical region in $\mathbf{Z}^{N}$ with $N:=n+1$; this ensures a coverage probability of at least $1-\epsilon$, where $\epsilon$ is the significance level used in $A$. Under randomness, we can produce non-vacuous (i.e., different from $\mathbf{Z}$) prediction sets at significance level $3\sqrt{N}\mathrm{e}^{-N}$ even in situations where no non-vacuous prediction sets are possible under exchangeability at any non-trivial significance level $\epsilon$ (i.e., at any $\epsilon<1$). It is instructive to compare this with conformal prediction, where non-vacuous prediction sets are only possible at much larger significance levels of at least $1/N$ [27, Sect. 11.4.4].
@@ 3.2 Kolmogorov’s and Martin-Löf’s work on Bernoulli sequences
The assumptions of randomness and exchangeability, in different guises, have also played important roles in the foundations of frequentist statistics. The standard measure-theoretic foundations of probability were put forward in Kolmogorov’s Grundbegriffe [13], but Kolmogorov did not believe that they were sufficient for applications of probability. As he pointed out in the Grundbegriffe [13, footnote 4], in his frequentist analysis of the applications of probability he was following Richard von Mises. However, a big difference between Kolmogorov’s and von Mises’s approaches was that von Mises’s was based on infinite sequences, whereas Kolmogorov believed that infinite sequences, being empirically non-existent, had no place in discussions of real-world applications of probability. Interestingly, because of this he even objected against publication in Russian Mathematical Surveys (a journal that he edited at the time) of a planned paper about infinite random sequences by Uspensky, Shen, and Semenov; see Kolmogorov’s letter to Uspensky of June 1983 cited in [23, note 14].
@@
For a long time Kolmogorov believed that no frequentist concept of probability can be developed for finite sequences [14, Sect. 1], but in 1963 he published his first attempt in this direction [14, Sect. 2]. The attempt, however, was ‘‘incomplete’’ (as he characterizes it in [15, Sect. 4]), and he greatly improved on it in 1968 [16, Sect. 2] (this paper is based on his 1967 talk). It appears that the details of Kolmogorov’s improved approach first appeared in print in Martin-Löf’s 1966 paper [20, Sect. 5].
@@
Both Kolmogorov and Martin-Löf consider binary sequences and define what they call Bernoulli sequences, i.e., sequences that can be plausibly obtained as result of IID observations. While their informal explanations clearly show that they are interested in the assumption of randomness, [^1] This is a relevant quote from Kolmogorov [16, Sect. 2]: “let us consider how we imagine a sequence of zeros and ones appearing as the result of independent trials with probability $p$ of obtaining a one at each trial.” their formal definitions are about the assumption of exchangeability. Let me give essentially Martin-Löf’s definition; I will use slightly different terminology, but my definition will be equivalent to Martin-Löf’s. This definition will rely on some basic notions of the theory of algorithms, but it will not be used outside this subsection, and the reader can skip the rest of the subsection without interrupting the flow of ideas.
@@
Let us define exchangeability p-variables and randomness p-variables as in the previous section, by (6) and (7), but replacing $\mathbf{Z}^{\infty}$ with $\mathbf{Z}^{N}$. We consider families $P_{N}$, $N\in\{1,2,\dots\}$, of exchangeability p-variables such that $P_{N}(x)$ is upper semicomputable as function of $N$ and $x\in\mathbf{Z}^{N}$ (where the upper semicomputability means that, for a computable function $f$ taking rational values, $P_{N}(x)=\inf_{k}f(N,x,k)$, $k$ ranging over the natural numbers). There exists a smallest, to within a constant factor, function $(N,x)\mapsto P_{N}(x)$ of this kind, and Martin-Löf defines
@@
Another equivalent definition of $m$ is given by Kolmogorov in [16, Sect. 2] in terms of his notion of complexity (and Martin-Löf proves the equivalence in [20, Sect. 5]). A binary sequence $x$ is called a Bernoulli sequence if $m(x)$ is small; this is an informal notion, but we can prove mathematical results about the function $m$ (albeit only with the $O(1)$ accuracy).
Replacing the assumption of exchangeability by that of randomness, we get a function that we denote by $\operatorname{\mathit{m}^{\mathrm{R}}}$ instead of $m$. (A natural notation for $m$ in view of our notations $\operatorname{\mathbb{P}^{\mathrm{X}}}$ and $\operatorname{\mathbb{E}^{\mathrm{X}}}$ would be $\operatorname{\mathit{m}^{\mathrm{X}}}$, but $m$ is what Martin-Löf used in [20, Sect. 5].)
@@
the inequality holding from some $N$ on; this is stated without proof in [26, Theorem 4]. We can see that the difference between $m$ and $\operatorname{\mathit{m}^{\mathrm{R}}}$ is very substantial.
@@ 3.3 Versions of de Finetti’s theorem for finite sequences of observations
An instructive finite form of de Finetti’s theorem was derived by Diaconis and Freedman [7] (with an early version given already at the very end of [11]). Since the assumptions of exchangeability and randomness are so different for a fixed length $N$, we have to consider different lengths for imposing exchangeability and for claiming randomness. In the abstract of [6], which states the Diaconis–Freedman result for binary sequences, Diaconis summarizes this result thus: ‘‘an exchangeable sequence of length $r$ which can be extended to an exchangeable sequence of length $k$ is almost a mixture of independent experiments, the error going to zero like $1/k$’’. It is essential that $k$ and $r$ should be different here, ideally $k\gg r$.
@@ 4.1 The case of infinite $\mathbf{Z}$
and $S_{N}$ stands for the symmetric group of all permutations of $\{1,\dots,N\}$ (this follows from, e.g., [27, Lemma A.3]). Therefore, it suffices to consider only $F$ that are non-zero on one orbit only. Let $F$ be non-zero on the orbit generated by a sequence $z_{1},\dots,z_{N}$ in $\mathbf{Z}^{N}$ containing $K$ distinct elements of $\mathbf{Z}$ with multiplicities $n_{1},\dots,n_{K}$ (so that $n_{1}+\dots+n_{K}=N$). The largest product probability $Q^{N}$ of an element of this orbit is
@@ 4.2 Smaller observation spaces $\mathbf{Z}$
The inequality (14) in Proposition 4.1 is tight in the case of an infinite $\mathbf{Z}$, which we relaxed in the statement of the proposition to what is actually used in the proof, $\left|\mathbf{Z}\right|\geq N$. The following proposition extends this to smaller $\mathbf{Z}$.
Proposition 4.2.
Suppose $K:=\left|\mathbf{Z}\right|<N$. For any random variable $F:\mathbf{Z}^{N}\to[0,\infty)$, $\operatorname{\mathbb{E}^{\mathrm{X}}}(F)\leq C\operatorname{\mathbb{E}^{\mathrm{R}}}(F)$, where
@@
Without loss of generality, we set $\mathbf{Z}:=\{1,\dots,K\}$ and proceed as in the proof of Proposition 4.1. Let $n_{k}$ be the number of times that $k\in\mathbf{Z}$ occurs in $z_{1},\dots,z_{N}$ (so that now $n_{k}=0$ is possible, in which case $n_{k}^{n_{k}}:=1$ and $n_{k}!=1$). Now instead of the inequality $\geq$’’ in (15) we use the convexity of the function $\log(n^{n}/n!)=n\log n-\log(n!)$ in $n$ (see Lemma 4.3 below). If the split $n_{k}$, $k=1,\dots,K$, of $N$ is not balanced, we can find $n_{k_{1}}$ and $n_{k_{2}}$ such that $n_{k_{1}}-n_{k_{2}}\geq 2$. By the convexity, the penultimate expression in the chain (15) cannot increase if we move $n_{k_{1}}$ and $n_{k_{2}}$ towards each other by redefining $n_{k_{1}}:=n_{k_{1}}-1$ and $n_{k_{2}}:=n_{k_{2}}+1$. Repeating this operation we arrive at a balanced split.
@@
The following lemma was used in the proof of Proposition 4.2.
Lemma 4.3.
@@
Figure 1: The graphs of $\log_{10}C$ for $N=10^{3}$ (thick red line, with some values given on the left) and $N=10^{6}$ (thin blue line, with values on the right).
Figure 1 shows $\log_{10}C$ as a function of $K\in\{1,\dots,N\}$ for two values of $N$, $10^{3}$ (red, with the scale of $\log_{10}C$ on the left) and $10^{6}$ (blue, with the scale on the right). Both graphs become horizontal to the right of $K:=N$ (not shown). We can see that the shapes of the two graphs are very similar. The blue line (for $10^{6}$) was drawn after the red one (for $10^{3}$), and the latter is thicker in order to be able to see the small difference between the graphs. When implementing the formula (16) for $C$, it is convenient to compute the number of $n_{k}=\lceil N/K\rceil$ in (17) as $N\bmod K$.
The vertical dotted lines in Fig. 1 are drawn through the observation $N/2$ (the right line) and the nearest observation to $N/3$ (the left one). The quantitative behaviour of the red and blue lines, namely the strictly exponential growth (linear on the log scale of Fig. 1) of both graphs between, roughly, $N/2$ and $N$, $N/3$ and $N/2$, etc., is easy to understand. When $K$ increases by 1 between $N/2$ and $N$, one of the $n_{k}=2$ in (17) gets replaced by two $n_{k}$, namely 1 and 1. Therefore, the expression for $C$ given by (16) gets multiplied by $2^{2}/2!=2$, and so the slope of the red and blue lines between $N/2$ and $N$ in Fig. 1 is $\log_{10}2\approx 0.301$. When $K$ increases by 1 between $N/3$ and $N/2$, a block $(3,3)$ of two $n_{k}$ in (17) gets replaced by a block $(2,2,2)$ of three $n_{k}$. The expression for $C$ given by (16) gets multiplied by
@@
and so the slope of the red and blue lines immediately to the left of $N/3$ in Fig. 1 is $\log_{10}(2^{19}/3^{11})\approx 0.471$. The slope keeps increasing as we move right to left.
One of the interesting cases for a finite $\mathbf{Z}$ is where $K:=\left|\mathbf{Z}\right|$ is fixed while the number of observations $N$ varies. (This corresponds to the bottom left corner of Fig. 1, where the slope becomes infinite ‘‘under microscope’’.) Applying Stirling’s formula to (16) we then obtain
$C=\Theta(N^{(K-1)/2}).$

### v2

@@ Abstract
Randomness (in the sense of being generated in an IID fashion) and exchangeability are standard assumptions in nonparametric statistics and machine learning, and relations between them have been a popular topic of research. This short paper draws the reader’s attention to the fact that, while for infinite sequences of observations the two assumptions are almost indistinguishable, the difference between them becomes very significant for finite sequences of a given length.
The version of this paper at http://alrw.net (Working Paper 45) is updated most often.
@@ 1 Introduction
In this paper we will discuss two important assumptions about sequences of observations, exchangeability and randomness, using the word ‘‘randomness’’ in a somewhat old-fashioned sense of the individual observations being independent and identically distributed (following [20, Chap. 7], which used the standard terminology for its time, and [28]).
The relationship between exchangeability and randomness is very different in the cases of infinite sequences and finite sequences. In the former case, there is hardly any difference between the two assumptions. But in the latter, the difference may be vast. The study of this relationship has a long history, which will also be briefly reviewed.
@@
According to Kolmogorov, and it is difficult to argue with his point of view, infinite sequences are empirically vacuous; we never observe them in reality. Therefore, he always insisted on either studying finite sequences or at least keeping them in mind. In fact, de Finetti’s results can also be stated in terms of finite sequences, provided we consider sequences of different lengths. In this paper we will briefly discuss two modes of statistical hypothesis testing involving finite sequences, online and batch; the former corresponds to de Finetti’s setting and the latter to Kolmogorov’s.
@@
This paper has little novelty, and all of its mathematical results are very simple. But the reader might find the scale of the difference between exchangeability and randomness for infinite observation spaces surprising.
@@ 2 De Finetti’s theorem
Suppose we observe a data sequence $z_{1},z_{2},\dots\in\mathbf{Z}$ consisting of observations $z_{n}\in\mathbf{Z}$ that are elements of some measurable space $\mathbf{Z}$, the observation space. The assumption of randomness is standard in machine learning: the data is coming from the product probability measure $Q^{\infty}$ for some $Q\in\mathfrak{P}(\mathbf{Z})$, where $\mathfrak{P}(Z)$ stands for the measurable space of all probability measures on a measurable space $Z$ (with the $\sigma$-algebra generated by the mappings $Q\mapsto Q(A)$, $A$ being an event in $Z$).
Remark 1.
Effects of various distribution shifts have also been widely studied, but in this paper we concentrate on the basic IID case. In machine learning we often have observations $z=(x,y)$ consisting of an object $x$ and its label $y$, but we do not insist on this.
@@
By de Finetti’s classical theorem each exchangeable probability measure $R$ on $\mathbb{R}^{\infty}$ is a mixture of product distributions:
@@
for some $\mu\in\mathfrak{P}(\mathfrak{P}(\mathbb{R}))$. This was established by de Finetti [4, Chap. 4]. Hewitt and Savage [11, Theorem 7.3 and its discussion later in Sect. 7] note that we can trivially replace $\mathbb{R}$ by any Borel space $\mathbf{Z}$ and point out that, to the best of their knowledge, every measurable space known to have importance in applied science is Borel. For example, every Polish space (complete separable metric space with its Borel $\sigma$-algebra) is Borel. On the other hand, there exists a separable metric space with its Borel $\sigma$-algebra for which (1) can be violated [8]. Let us assume in the rest of this paper that the observation space $\mathbf{Z}$ is Borel.
It is well known that de Finetti’s theorem fails if we simply replace $\infty$ by a finite $N$ in (1); see, e.g., [3, Sect. 4.7.1] (and [6, Sect. 1] for a further discussion of the extent to which de Finetti’s theorem can fail for $N=2$ and $\mathbf{Z}=\{0,1\}$).
@@
$\Gamma(z_{1},\dots,z_{n}):=\bigl\{(z_{n+1},z_{n+2},\dots):(z_{1},\dots,z_{n},z_{n+1},z_{n+2},\dots)\notin A\bigr\}$ (5)
for the future observations after observing $z_{1},\dots,z_{n}$, where $A$ is a critical region at some significance level $\epsilon$. Under both exchangeability and randomness, the coverage probability of the prediction set (5) will be at least $1-\epsilon$.
There are two popular generalizations of critical regions, p-variables and e-variables, and both also produce identical results under exchangeability and randomness. Let us first check this for p-variables. According to the general definition (see, e.g., [22, Definition 1.2]), an exchangeability p-variable is a random variable $P:\mathbf{Z}^{\infty}\to[0,1]$ such that, for all $\epsilon\in(0,1)$,
@@
Since $\operatorname{\mathbb{P}^{\mathrm{X}}}=\operatorname{\mathbb{P}^{\mathrm{R}}}$, the classes of exchangeability and randomness p-variables coincide.
Again according to the general definition [22, Definition 1.2] (see also [10, (1)]), an exchangeability e-variable is a measurable function $F:\mathbf{Z}^{\infty}\to[0,\infty]$ such that $\operatorname{\mathbb{E}^{\mathrm{X}}}(F)\leq 1$, where
@@
3 Finite sequences of observations
@@ 3.1 Batch setting
$\frac{N}{N}\frac{N-1}{N}\dots\frac{1}{N}=\frac{N!}{N^{N}}\sim\sqrt{2\pi N}\mathrm{e}^{-N};$ (11)
@@
Remark 2.
A convenient, albeit asymptotically much cruder, version of the inequality (12) is
$\operatorname{\mathbb{P}^{\mathrm{R}}}(A)\leq 2^{-N+1}\leq 1=\operatorname{\mathbb{P}^{\mathrm{X}}}(A).$
Remark 3.
@@
Similarly to (5), we can invert exchangeability and randomness critical regions in the batch mode and use them for one-step-ahead prediction (whereas prediction in (5) was infinitely many steps ahead). Given an observed sequence $z_{1},\dots,z_{n}$, we output
@@
as our prediction set for the next observation, where $A$ is a critical region in $\mathbf{Z}^{N}$ with $N:=n+1$; this ensures a coverage probability of at least $1-\epsilon$, where $\epsilon$ is the significance level used in $A$. Under randomness, we can produce non-vacuous (i.e., different from $\mathbf{Z}$) prediction sets at significance level $3\sqrt{N}\mathrm{e}^{-N}$ even in situations where no non-vacuous prediction sets are possible under exchangeability at any non-trivial significance level $\epsilon$ (i.e., at any $\epsilon<1$). It is instructive to compare this with conformal prediction, where non-vacuous prediction sets are only possible at much larger significance levels of at least $1/N$ [28, Sect. 11.4.4].
@@ 3.2 Kolmogorov’s and Martin-Löf’s work on Bernoulli sequences
The assumptions of randomness and exchangeability, in different guises, have also played important roles in the foundations of frequentist statistics. The standard measure-theoretic foundations of probability were put forward in Kolmogorov’s Grundbegriffe [14], but Kolmogorov did not believe that they were sufficient for applications of probability. As he pointed out in the Grundbegriffe [14, footnote 4], in his frequentist analysis of the applications of probability he was following Richard von Mises. However, a big difference between Kolmogorov’s and von Mises’s approaches was that von Mises’s was based on infinite sequences, whereas Kolmogorov believed that infinite sequences, being empirically non-existent, had no place in discussions of real-world applications of probability. Interestingly, because of this he even objected against publication in Russian Mathematical Surveys (a journal that he edited at the time) of a planned paper about infinite random sequences by his student Uspensky and close collaborators Shen and Semenov; see Kolmogorov’s letter to Uspensky of June 1983 cited in [24, note 14].
@@
For a long time Kolmogorov believed that no frequentist concept of probability can be developed for finite sequences [15, Sect. 1], but in 1963 he published his first attempt in this direction [15, Sect. 2]. The attempt, however, was ‘‘incomplete’’ (as he characterizes it in [16, Sect. 4]), and he greatly improved on it in 1968 [17, Sect. 2] (this paper is based on his 1967 talk). It appears that the details of Kolmogorov’s improved approach first appeared in print in Martin-Löf’s 1966 paper [21, Sect. 5].
@@
Both Kolmogorov and Martin-Löf consider binary sequences and define what they call Bernoulli sequences, i.e., sequences that can be plausibly obtained as result of IID observations. While their informal explanations clearly show that they are interested in the assumption of randomness, [^1] This is a relevant quote from Kolmogorov [17, Sect. 2]: “let us consider how we imagine a sequence of zeros and ones appearing as the result of independent trials with probability $p$ of obtaining a one at each trial.” their formal definitions are about the assumption of exchangeability. Let me give essentially Martin-Löf’s definition; I will use slightly different terminology, but my definition will be equivalent to Martin-Löf’s. This definition will rely on some basic notions of the theory of algorithms, but it will not be used outside this subsection, and the reader can skip the rest of the subsection without interrupting the flow of ideas.
@@
Let us define exchangeability p-variables and randomness p-variables as in the previous section, by (6) and (7), but replacing $\mathbf{Z}^{\infty}$ with $\mathbf{Z}^{N}$ for a finite $N$. We consider families $P_{N}$, $N\in\{1,2,\dots\}$, of exchangeability p-variables such that $P_{N}(x)$ is upper semicomputable as function of $N$ and $x\in\mathbf{Z}^{N}$ (where the upper semicomputability means that, for a computable function $f$ taking rational values, $P_{N}(x)=\inf_{k}f(N,x,k)$, $k$ ranging over the natural numbers). There exists a smallest, to within a constant factor, function $(N,x)\mapsto P_{N}(x)$ of this kind, and Martin-Löf defines
@@
Another equivalent definition of $m$ is given by Kolmogorov in [17, Sect. 2] in terms of his notion of complexity (and Martin-Löf proves the equivalence in [21, Sect. 5]). A binary sequence $x$ is called a Bernoulli sequence if $m(x)$ is small; this is an informal notion, but we can prove mathematical results about the function $m$ (albeit only with the $O(1)$ accuracy).
Replacing the assumption of exchangeability by that of randomness, we get a function that we denote by $\operatorname{\mathit{m}^{\mathrm{R}}}$ instead of $m$. (A natural notation for $m$ in view of our notations $\operatorname{\mathbb{P}^{\mathrm{X}}}$ and $\operatorname{\mathbb{E}^{\mathrm{X}}}$ would have been $\operatorname{\mathit{m}^{\mathrm{X}}}$, but $m$ is what Martin-Löf used in [21, Sect. 5].)
@@
the inequality holding from some $N$ on; this is stated without proof in [27, Theorem 4]. We can see that the difference between $m$ and $\operatorname{\mathit{m}^{\mathrm{R}}}$ is very substantial.
@@ 3.3 Versions of de Finetti’s theorem for finite sequences of observations
An instructive finite form of de Finetti’s theorem was derived by Diaconis and Freedman [7] (with an early version given already at the very end of [11]). Since the assumptions of exchangeability and randomness are so different for a fixed length $N$, we have to consider different lengths for imposing exchangeability and for claiming randomness. In the abstract of [6], which states the Diaconis–Freedman result for binary sequences, Diaconis summarizes this result thus: ‘‘an exchangeable sequence of length $r$ which can be extended to an exchangeable sequence of length $k$ is almost a mixture of independent experiments, the error going to zero like $1/k$’’. It is essential that $k$ and $r$ should be different here, ideally $k\gg r$. There have been several recent information-theoretic developments of this idea; see [12, Corollary 1] for a particularly strong result.
@@ 4.1 The case of infinite $\mathbf{Z}$
and $S_{N}$ stands for the symmetric group of all permutations of $\{1,\dots,N\}$ (this follows from, e.g., [28, Lemma A.3]). Therefore, it suffices to consider only $F$ that are non-zero on one orbit only. Let $F$ be non-zero on the orbit generated by a sequence $z_{1},\dots,z_{N}$ in $\mathbf{Z}^{N}$ containing $K$ distinct elements of $\mathbf{Z}$ with multiplicities $n_{1},\dots,n_{K}$ (so that $n_{1}+\dots+n_{K}=N$). The largest product probability $Q^{N}$ of an element of this orbit is
@@ 4.2 Smaller observation spaces $\mathbf{Z}$
The inequality (14) in Proposition 4 is tight in the case of an infinite $\mathbf{Z}$, which we relaxed in the statement of the proposition to what is actually used in the proof, $\left|\mathbf{Z}\right|\geq N$. The following proposition extends (14) to smaller $\mathbf{Z}$.
Proposition 5.
Suppose $K:=\left|\mathbf{Z}\right|<N$. For any random variable $F:\mathbf{Z}^{N}\to[0,\infty)$, we have $\operatorname{\mathbb{E}^{\mathrm{X}}}(F)\leq C\operatorname{\mathbb{E}^{\mathrm{R}}}(F)$, where
@@
Without loss of generality, we set $\mathbf{Z}:=\{1,\dots,K\}$ and proceed as in the proof of Proposition 4. Let $n_{k}$ be the number of times that $k\in\mathbf{Z}$ occurs in $z_{1},\dots,z_{N}$ (so that now $n_{k}=0$ is possible, in which case $n_{k}^{n_{k}}:=1$ and $n_{k}!=1$). Now instead of the inequality $\geq$’’ in (15) we use the convexity of the function $\log(n^{n}/n!)=n\log n-\log(n!)$ in $n$ (see Lemma 6 below). If the split $n_{k}$, $k=1,\dots,K$, of $N$ is not balanced, we can find $n_{k_{1}}$ and $n_{k_{2}}$ such that $n_{k_{1}}-n_{k_{2}}\geq 2$. By the convexity, the penultimate expression in the chain (15) cannot increase if we move $n_{k_{1}}$ and $n_{k_{2}}$ towards each other by redefining $n_{k_{1}}:=n_{k_{1}}-1$ and $n_{k_{2}}:=n_{k_{2}}+1$. Repeating this operation we arrive at a balanced split.
@@
The following lemma was used in the proof of Proposition 5.
Lemma 6.
@@
Figure 1: The graphs of $\log_{10}C$ for $N=10^{3}$ (thick red line, with some values given on the left) and $N=10^{6}$ (thin blue line, with values on the right). The yellow line represents the approximation described in Remark 7.
Figure 1 shows $\log_{10}C$ as a function of $K\in\{1,\dots,N\}$ for two values of $N$, $10^{3}$ (red, with the scale of $\log_{10}C$ on the left) and $10^{6}$ (blue, with the scale on the right). Both graphs become horizontal to the right of $K:=N$ (not shown in Fig. 1). We can see that the shapes of the two graphs are very similar. The blue line (for $10^{6}$) was drawn after the red one (for $10^{3}$), and the latter is thicker in order to be able to see the small difference between the graphs. When implementing the formula (16) for $C$, it is convenient to compute the number of $n_{k}=\lceil N/K\rceil$ in (17) as $N\bmod K$.
The vertical dotted lines in Fig. 1 are drawn through the observation $N/2$ (the right line) and the nearest observation to $N/3$ (the left one). The qualitative behaviour of the red and blue lines, namely the strictly exponential growth (linear on the log scale of Fig. 1) of both graphs between, roughly, $N/2$ and $N$, $N/3$ and $N/2$, etc., is easy to understand. When $K$ increases by 1 between $N/2$ and $N$, one of the $n_{k}=2$ in (17) gets replaced by two $n_{k}$, namely 1 and 1. Therefore, the expression for $C$ given by (16) gets multiplied by $2^{2}/2!=2$, and so the slope of the red and blue lines between $N/2$ and $N$ in Fig. 1 is $\log_{10}2\approx 0.301$. When $K$ increases by 1 between $N/3$ and $N/2$, a block $(3,3)$ of two $n_{k}$ in (17) gets replaced by a block $(2,2,2)$ of three $n_{k}$. The expression for $C$ given by (16) gets multiplied by
@@
and so the slope of the red and blue lines immediately to the left of $N/3$ in Fig. 1 is $\log_{10}(2^{19}/3^{11})\approx 0.471$. The slope keeps increasing as we move left.
Remark 7.
Let us replace the decimal logarithms $\log_{10}$ by natural $\ln$ in the graphs shown in Fig. 1 for $N\in\{10^{3},10^{6}\}$. This will not change the shape of the graphs and will only change the labels on the axes; the upper limit of the range of $\log C$ will now become close to $N$ (this will be checked in the appendix). An interesting function is the limit $L$ of these graphs as $N\to\infty$ with both axes rescaled by dividing by $N$ (so that the slopes remain unchanged). It can be defined as the continuous piecewise-linear function $L:[0,1]\to[0,\infty)$ satisfying $L(0):=0$ and
$L^{\prime}(x):=\ln\left(\left(\frac{n!}{n^{n}}\right)^{n+1}\left(\frac{(n+1)!}{(n+1)^{n+1}}\right)^{-n}\right)\text{ for all }x\in\left(\frac{1}{n+1},\frac{1}{n}\right)$ (19)
and for all $n\in\{1,2,\dots\}$. As $x\to 0$,
$L(x)=-\frac{1}{2}x\ln x+\frac{\ln(2\pi)}{2}x+o(x),$
and the right-hand side without the ‘‘+o(x){}+o(x)’’ and with $\log_{10}$ in place of $\ln$ is shown as the yellow line in Fig. 1. The final value of the approximation $\frac{1}{2}x\ln x+\frac{\ln(2\pi)}{2}x$ at $x=1$ is approximately 0.919, which is not so different from $L(1)=1$. We can see that the slope of $L(x)$ is infinity at $x=0$.
Another interesting case for a finite $\mathbf{Z}$ is where $K:=\left|\mathbf{Z}\right|$ is fixed while the number of observations $N$ varies. Applying Stirling’s formula to (16) we then obtain
$C=\Theta(N^{(K-1)/2}).$ (20)